OpenStudy (anonymous):

Solve for y: 2/3y + 4 = 6 Answer y = 3 y = -7 y = 7 y = -3

jimthompson5910 (jim_thompson5910):

Any ideas on the first step?

OpenStudy (anonymous):

-4 on both sides?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 2/3y = 2

jimthompson5910 (jim_thompson5910):

what's next

mathslover (mathslover):

\[\huge{\frac{2y}{3}+4=6}\] \[\huge{\frac{2y}{3}=2}\] \[\huge{2y=6}\] \[\huge{y=\frac{6}{2}=3}\]

OpenStudy (anonymous):

I thought it equaled 2 also, but I don't have 2 as an option D: D: D:

jimthompson5910 (jim_thompson5910):

You want to move that 3 over, so multiply both sides by 3 2/3y = 2 (2/3y)*3 = 2*3 2y = 6

jimthompson5910 (jim_thompson5910):

Then what's next?